a) Find an equation of the tangent line at the origin to the curve .


1+sin(x+y)=xy+cosy


1
Expert's answer
2022-01-10T16:13:41-0500

tangent line:

"y-y_0=y'(x_0)(x-x_0)"


we have:

"(x_0,y_0)=(0,0)"

"y'cos(x+y)=y+xy'-y'siny"


"y'=\frac{y}{cos(x+y)+siny-x}"


"y'(0)=0"


then, equation of tangent line:

"y=0"